3. Consider an economy that initially has a labor force of 4000 workers. Of these workers, 3000 are employed . The economy enters a recession. The total number of employed workers falls by 10%, and in addition 1% of the labour force become discouraged at the prospect of finding a job and stop searching for a job for more than 4 weeks Can you calculate the employment rate given the information above? Explain. Calculate the unemployment rate before, and after, the recession. Are you able to calculate the labour force participation rate given the information above? Explain. a) b) c) d) What is the unemployment rate accounting for discouraged workers? e) Using Okuns law, as given in class, calculate the expected percentage change in real GDP due to the recession.
Answer #1

B) Before recession

Labour force = 4000

Employed = 3000

Labour force = employed + unemployed

4000 = 3000+ unemployed

Unemployed = 1000

Unemployment rate =(1000/4000)*100 = 25%

Employment rate = (3000/4000)*100 = 75%

After recession

Labour force = 4000(1-1%) = (1-0.01) *4000

New Labour force = 3960

New employed = 3000(1-10%) = 3000(1-0.1)

New employed = 2700

Unemployed = 3960-2700 = 1260

Unemployment rate = (1260/3960)*100 = 31.82%

a) Employment rate after recession

Employment rate = (2700/3960)*100 = 68.18%

C) No I cannot calculate the labour force participation rate as the total adult population is not known.

Labour force participation rate = Labour force/ adult population

e)

Okun law

(Unemployment rate after recession - unemployment rate before recession)*2.5

(31.82-25)*2.5 = 17.05%

Expected % change in RGDP is 17.05%